# DirDist: A Metric for Comparing 3D Shapes Using Directional Distance Fields

### v4

@@ Abstract
Qualifying the discrepancy between 3D geometric models, which could be represented with either point clouds or triangle meshes, is a pivotal issue with board applications. Existing methods mainly focus on directly establishing the correspondence between two models and then aggregating point-wise distance between corresponding points, resulting in them being either inefficient or ineffective. In this paper, we propose DirDist, an efficient, effective, robust, and differentiable distance metric for 3D geometry data. Specifically, we construct DirDist based on the proposed implicit representation of 3D models, namely directional distance field (DDF), which defines the directional distances of 3D points to a model to capture its local surface geometry. We then transfer the discrepancy between two 3D geometric models as the discrepancy between their DDFs defined on an identical domain, naturally establishing model correspondence. To demonstrate the advantage of our DirDist, we explore various distance metric-driven 3D geometric modeling tasks, including template surface fitting, rigid registration, non-rigid registration, scene flow estimation and human pose optimization. Extensive experiments show that our DirDist achieves significantly higher accuracy under all tasks. As a generic distance metric, DirDist has the potential to advance the field of 3D geometric modeling. The source code is available at https://github.com/rsy6318/DirDist.
@@ I Introduction
To address this fundamental and challenging issue of quantifying the discrepancy between 3D geometric models, we propose a novel robust, efficient, and effective distance metric called DirDist.Distinguished from existing metrics, DirDist emphasizes the implicit representation of 3D geometric models, as depicted in Fig. 1(e). Methodologically, for any pair of 3D models, we begin by constructing their individual implicit fields, referred to as directional distance fields (DDFs). The discrepancy between these fields serves as an indicator of the distance between the corresponding 3D geometric models. To compute DirDist, we generate a set of reference points that are distributed near the models and shared by both models. For each reference point, we calculate its directional distances to the two models, capturing the local surface geometry. By taking the weighted average of the directional distance discrepancies contributed by all reference points, we obtain the final DirDist. The use of DDFs enables DirDist to robustly handle optimization convergence issues by effectively capturing surface geometry. Moreover, DirDist does not require a direct correspondence establishment process, resulting in high efficiency.
Notably, DirDist is differentiable, allowing seamless integration as a module in various tasks. We leverage DirDist to develop methods for diverse 3D geometric modeling and processing tasks, including template surface fitting, rigid and non-rigid registration, scene flow estimation, and human pose optimization. Extensive experiments validate the superiority of DirDist in terms of robustness, efficiency, and effectiveness
@@
  • an efficient, effective, robust, and generic distance metric for 3D geometry data, dubbed DirDist;
  • various state-of-the-art geometric modeling and processing methods driven by our DirDist;
The remainder of this paper is structured as follows. Section II provides a brief overview of existing literature on 3D geometry representations, distance metrics for 3D geometry data, and distance metric-driven 3D geometry processing. Section III introduces our proposed DirDist in detail. Section IV outlines the specifics of the five general distance metric-driven 3D geometric tasks, followed by comprehensive experiments on these tasks in Section V. Finally, Section VI concludes this work.
@@ III-B Problem Statement and Motivation
In what follows, we will detail the generation of reference points in Sec. III-C and the definition of DDF in Sec. III-D, leading to our efficient yet effective distance metric for 3D geometry data named DirDist in Sec. III-E.
@@ III-E DDF-based Distance Metric
Based on the previously introduced DDF, we can reformulate the representations of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ as $\mathcal{S}_{1}=\{\mathcal{F}_{\mathcal{S}_{1}}(\mathbf{q})|\mathbf{q}\in\mathbf{Q}\}$ and $\mathcal{S}_{2}=\{\mathcal{F}_{\mathcal{S}_{2}}(\mathbf{q})|\mathbf{q}\in\mathbf{Q}\}$, respectively, where the correspondence between $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ is established in a indirect fashion through $\mathbf{Q}$. We finally define the our distance metric for 3D geometry data named DirDist as
$\mathcal{D}_{\rm DirDist}(\mathcal{S}_{1},\mathcal{S}_{2})=\iiint_{\mathbf{Q}}s(\mathbf{q})\cdot d(\mathbf{q},\mathcal{S}_{1},\mathcal{S}_{2}){\rm d}\mathbf{q},$ (7)
@@
The proposed DirDist in Eq. (7) possesses essential properties that align with standard distance metrics, including non-negativity, symmetry, the identity of indiscernibles, and the triangle inequality. Obviously, the first three properties are satisfied once the reference points are given. Theorem 1 shows that it also satisfies the triangle inequality. In addition, under some specific settings, existing distance metrics, such as CD and P2F, can be regarded as special cases of our DirDist. Theorems 2 and 3 illustrate these properties.
@@
Given two point clouds, $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$, when the reference points are defined as $\mathbf{Q}=\mathbf{P}_{1}\cup\mathbf{P}_{2}$, and the DDF is defined as $\mathcal{F}=f$ with parameters $K=1$ and $\beta=0$, DirDist becomes equivalent to CD.
@@
where NN is the 1-NN operation. Thus, DirDist under these reference points is
@@
Similarity, DirDist under the reference points $\mathbf{q}\in\mathbf{P}_{2}$ is
@@
Obviously, $\mathcal{D}_{\rm DirDist}=\mathcal{D}_{1}+\mathcal{D}_{2}$ is the same as CD.
@@
Given two triangle meshes, $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, when the reference points $\mathbf{Q}=\{\mathbf{q}_{i}\}_{i=1}^{M}$ are sampled on the surface, and the DDF is defined as $\mathcal{F}=f$ with parameter $\beta=0$, DirDist becomes equivalent to the P2F distance.
@@
where NP is the nearest point of the given point on the surface. Thus, DirDist unde these reference points is
@@
Similarity, DirDist under the reference points sampled from $\mathcal{S}_{2}$, $\mathbf{Q}_{2}=\{\mathbf{q}_{i}\}_{i=1}^{M_{2}}$ is
@@
Obviously, $\mathcal{D}_{\rm DirDist}=\mathcal{D}_{1}+\mathcal{D}_{2}$ is the same as P2F distance. ∎
@@ IV Distance Metric-driven 3D Geometric Modeling
To demonstrate the superiority of the proposed DirDist, we apply it across an extensive array of fundamental 3D geometric modeling tasks, encompassing template-based surface reconstruction, rigid and non-rigid surface registration, scene flow estimation, and SMPL registration. In what follows, we will introduce the detailed method of each task.
@@ V-A2 Comparisons
We chose two baseline distance metrics for comparison, i.e., CD and P2F. For a fair comparison, the numbers of sampled points for computing CD and P2F were kept the same as our DirDist. Besides, we also made a comparison with MDA [53], where a one-sided CD is used as the distance metric and the other settings were kept the same as their original paper. We utilized Normal Consistency (NC) and F-Score with thresholds of 0.5% and 1%, denoted as F-0.005 and F-0.01, as the evaluation metrics. We refer the readers to [28] for the detailed definitions. During the evaluation, 500K points were sampled on the deformed and target surfaces. Table I and Fig. 8 show the numerical and visual results, respectively, where both quantitative accuracy and visual quality of deformed shapes with our DirDist are much better than baseline methods. While MDA exhibits superior visual results compared to CD and P2F, its numerical results about F-0.005 and F-0.01 in Table I is worse than both. This discrepancy arises from the fact that the deformed surfaces produced by MDA are excessively smoothed, leading to a convergence to local optima, because of its loss function. Although the overall shape is similar, this oversmoothing effect is evident in the error maps illustrated in Fig. 9. Table II presents the running time and GPU memory costs of different distance metrics per iteration. Notably, our DirDist incurs a computational cost similar to P2F but higher than CD. This is because both DirDist and P2F require identifying the closest points on triangle meshes, a more complex task compared to CD, which only involves finding the nearest points in point clouds.
@@ V-B2 Comparisons
We compared our DirDist with EMD [14], CD [15], DCD [17], and ARL [3]. Additionally, we made comparisons with two widely used registration methods: ICP [1] and SICP [48]. We employed two widely used evaluation metrics in the rigid registration task, Rotation Error (RE) and Translation Error (TE), to measure the registration accuracy, respectively defined as
@@
The numerical results are listed in Table III, where the refinement through our DirDist further improves the accruacy of MAC significantly, while after the refinement with other distance metric, the registration accuracy even decrease dynamically. Although the RE and TE of SICP are slightly better than ours, its SR is much lower than ours, showing its limitation. The visual results are shown in Fig. 10, and when dealing with the data with noise or outliers, the details of the point clouds become indistinct, our method still register the two point clouds successfully. To evaluate the performance of different methods insightly, we also used the Registration Recall (RR) under different thresholds of RE and TE to measure the registration accuracy. As shown in Fig. 11, the superiority of our DirDist is further verified. Additionally, Table IV presents the computational cost per iteration during the optimization process. As shown, EMD incurs the highest computational overhead due to the necessity of constructing a comprehensive bi-directional mapping between two point sets. Similarly, ARL requires significant computational resources, as it involves calculating the intersection of lines with the two point clouds. In contrast, CD, BCD, and our DirDist exhibit comparable computational costs. However, DirDist achieves superior registration results compared to these methods, demonstrating its effectiveness and efficiency in rigid point cloud registration.
@@ V-C2 Comparisons
We compared our DirDist with CD and the P2F distance. For a fair comparison, we kept the number of sampled points in CD and the P2F distance the same as the generated reference points of our DirDist. In addition, we also compared with the state-of-the-art optimization-based non-rigid registration method called AMM [4]. We computed RMSE of the estimated and ground truth vertices of the deformed source surfaces for quantitative evaluation. The results in Table V and Fig. 12 demonstrate the significant superiority of our DirDist. Additionally, we show the registration errors of all pairs in a sequence in Fig. 13 for a more comprehensive comparison, where it can be seen that the compared methods produce much larger RMSE values for pairs with significant motion, while our DirDist always works well. Additionally, Table VI presents the computational time and GPU memory requirements per iteration during the optimization process. Specifically, CD achieves the fastest running speed, as it only involves identifying the nearest points within the point clouds. In contrast, P2F and our proposed DirDist require more computational time due to the added complexity of locating the closest point on the triangle mesh, which is inherently more complex than finding the nearest points in a point cloud.
@@ V-D1 Implementation Details
We used the Flyingthings3D dataset [64], where $N_{\rm src}=N_{\rm tgt}=8192$. The hyperparameters involved in our DirDist were set as $K=5$ and $M=81920$. In order to adapt the density at each point’s location, the value of $\sigma$ was set as 3 times of the distance to its nearest point. For the optimization-based methods, we optimized the scene flow directly with the Adam optimizer for 500 iterations with a learning rate of 0.01. For unsupervised learning-based methods, we employed two state-of-the-art methods, namely NSFP [12] and SCOOP [13], both of which employ CD as the distance metric. We maintained the training settings identical to those specified in their original papers, with the only modification being the replacement of the alignment criterion with our DirDist.
@@ V-D2 Comparisons
Following [12, 13], we employed End Point Error (EPE), Flow Estimation Accuracy (Acc) with thresholds 0.05 and 0.1 (denoted as Acc-0.05 and Acc-0.1), and Outliers as the evaluation metrics. From Table VII and Fig. 14, it can be seen that our DirDist drives much more accurate scene flows than EMD, CD, and BCD under the optimization-based framework, and our DirDist further boosts the accuracy of SOTA unsupervised learning-based methods to a significant extent, demonstrating its superiority and the importance of the distance metric in 3D data modeling. Table VIII illustrates the computational cost of various distance metrics per iteration during the optimization process. Specifically, EMD demands significantly more computing resources due to the necessity of constructing a bi-directional mapping between two point sets. In contrast, CD, BCD, and our proposed DirDist exhibit comparable computational requirements. However, our DirDist outperforms the others in terms of flow estimation accuracy, highlighting its superior effectiveness and efficiency.
@@ V-E1 Implementation Details
We used the CAPE dataset [65], which contains numbers of clothed human models and their corresponding SMPL parameters. We selected two sequences from the whole dataset named ‘longlong basketball trial2’ and ‘blazerlong volleyball trial2’. In each sequence, we used the pose of the $i$-th frame as the initial parameter and the points sampled from the ($i$+5)-th frame as the simulated scanned point cloud, where the number of the sampled points was $2\times 10^{4}$. The hyperparameters invovled in our DirDist were set as $K=5$, $\beta=1.0$, and $M=6\times 10^{4}$. To bring the experimental setup closer to real-world conditions, where collected point clouds are typically incomplete, we positioned a virtual camera in front of the reference mesh and selectively sampled points exclusively from the visible faces, emulating the data collection process of an RGB-D camera. In such a case, the number of the sampled points was $10^{4}$ and the hyperparameters about our DirDist were kept the same. The optimization was conducted through Adam optimizer, running $10^{3}$ iterations with a learning rate of $5\times 10^{-4}$.

### v5

# DDM: A Metric for Comparing 3D Shapes Using Directional Distance Fields

@@ Abstract
Qualifying the discrepancy between 3D geometric models, which could be represented with either point clouds or triangle meshes, is a pivotal issue with board applications. Existing methods mainly focus on directly establishing the correspondence between two models and then aggregating point-wise distance between corresponding points, resulting in them being either inefficient or ineffective. In this paper, we propose DDM, an efficient, effective, robust, and differentiable distance metric for 3D geometry data. Specifically, we construct DDM based on the proposed implicit representation of 3D models, namely directional distance field (DDF), which defines the directional distances of 3D points to a model to capture its local surface geometry. We then transfer the discrepancy between two 3D geometric models as the discrepancy between their DDFs defined on an identical domain, naturally establishing model correspondence. To demonstrate the advantage of our DDM, we explore various distance metric-driven 3D geometric modeling tasks, including template surface fitting, rigid registration, non-rigid registration, scene flow estimation and human pose optimization. Extensive experiments show that our DDM achieves significantly higher accuracy under all tasks. As a generic distance metric, DDM has the potential to advance the field of 3D geometric modeling. The source code is available at https://github.com/rsy6318/DDM.
@@ I Introduction
To address this fundamental and challenging issue of quantifying the discrepancy between 3D geometric models, we propose a novel robust, efficient, and effective distance metric called DDM.Distinguished from existing metrics, DDM emphasizes the implicit representation of 3D geometric models, as depicted in Fig. 1(e). Methodologically, for any pair of 3D models, we begin by constructing their individual implicit fields, referred to as directional distance fields (DDFs). The discrepancy between these fields serves as an indicator of the distance between the corresponding 3D geometric models. To compute DDM, we generate a set of reference points that are distributed near the models and shared by both models. For each reference point, we calculate its directional distances to the two models, capturing the local surface geometry. By taking the weighted average of the directional distance discrepancies contributed by all reference points, we obtain the final DDM. The use of DDFs enables DDM to robustly handle optimization convergence issues by effectively capturing surface geometry. Moreover, DDM does not require a direct correspondence establishment process, resulting in high efficiency.
Notably, DDM is differentiable, allowing seamless integration as a module in various tasks. We leverage DDM to develop methods for diverse 3D geometric modeling and processing tasks, including template surface fitting, rigid and non-rigid registration, scene flow estimation, and human pose optimization. Extensive experiments validate the superiority of DDM in terms of robustness, efficiency, and effectiveness
@@
  • an efficient, effective, robust, and generic distance metric for 3D geometry data, dubbed DDM;
  • various state-of-the-art geometric modeling and processing methods driven by our DDM;
The remainder of this paper is structured as follows. Section II provides a brief overview of existing literature on 3D geometry representations, distance metrics for 3D geometry data, and distance metric-driven 3D geometry processing. Section III introduces our proposed DDM in detail. Section IV outlines the specifics of the five general distance metric-driven 3D geometric tasks, followed by comprehensive experiments on these tasks in Section V. Finally, Section VI concludes this work.
@@ III-B Problem Statement and Motivation
In what follows, we will detail the generation of reference points in Sec. III-C and the definition of DDF in Sec. III-D, leading to our efficient yet effective distance metric for 3D geometry data named DDM in Sec. III-E.
@@ III-E DDF-based Distance Metric
Based on the previously introduced DDF, we can reformulate the representations of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ as $\mathcal{S}_{1}=\{\mathcal{F}_{\mathcal{S}_{1}}(\mathbf{q})|\mathbf{q}\in\mathbf{Q}\}$ and $\mathcal{S}_{2}=\{\mathcal{F}_{\mathcal{S}_{2}}(\mathbf{q})|\mathbf{q}\in\mathbf{Q}\}$, respectively, where the correspondence between $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ is established in a indirect fashion through $\mathbf{Q}$. We finally define the our distance metric for 3D geometry data named DDM as
$\mathcal{D}_{\rm DDM}(\mathcal{S}_{1},\mathcal{S}_{2})=\iiint_{\mathbf{Q}}s(\mathbf{q})\cdot d(\mathbf{q},\mathcal{S}_{1},\mathcal{S}_{2}){\rm d}\mathbf{q},$ (7)
@@
The proposed DDM in Eq. (7) possesses essential properties that align with standard distance metrics, including non-negativity, symmetry, the identity of indiscernibles, and the triangle inequality. Obviously, the first three properties are satisfied once the reference points are given. Theorem 1 shows that it also satisfies the triangle inequality. In addition, under some specific settings, existing distance metrics, such as CD and P2F, can be regarded as special cases of our DDM. Theorems 2 and 3 illustrate these properties.
@@
Given two point clouds, $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$, when the reference points are defined as $\mathbf{Q}=\mathbf{P}_{1}\cup\mathbf{P}_{2}$, and the DDF is defined as $\mathcal{F}=f$ with parameters $K=1$ and $\beta=0$, DDM becomes equivalent to CD.
@@
where NN is the 1-NN operation. Thus, DDM under these reference points is
@@
Similarity, DDM under the reference points $\mathbf{q}\in\mathbf{P}_{2}$ is
@@
Obviously, $\mathcal{D}_{\rm DDM}=\mathcal{D}_{1}+\mathcal{D}_{2}$ is the same as CD.
@@
Given two triangle meshes, $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, when the reference points $\mathbf{Q}=\{\mathbf{q}_{i}\}_{i=1}^{M}$ are sampled on the surface, and the DDF is defined as $\mathcal{F}=f$ with parameter $\beta=0$, DDM becomes equivalent to the P2F distance.
@@
where NP is the nearest point of the given point on the surface. Thus, DDM unde these reference points is
@@
Similarity, DDM under the reference points sampled from $\mathcal{S}_{2}$, $\mathbf{Q}_{2}=\{\mathbf{q}_{i}\}_{i=1}^{M_{2}}$ is
@@
Obviously, $\mathcal{D}_{\rm DDM}=\mathcal{D}_{1}+\mathcal{D}_{2}$ is the same as P2F distance. ∎
@@ IV Distance Metric-driven 3D Geometric Modeling
To demonstrate the superiority of the proposed DDM, we apply it across an extensive array of fundamental 3D geometric modeling tasks, encompassing template-based surface reconstruction, rigid and non-rigid surface registration, scene flow estimation, and SMPL registration. In what follows, we will introduce the detailed method of each task.
@@ V-A2 Comparisons
We chose two baseline distance metrics for comparison, i.e., CD and P2F. For a fair comparison, the numbers of sampled points for computing CD and P2F were kept the same as our DDM. Besides, we also made a comparison with MDA [53], where a one-sided CD is used as the distance metric and the other settings were kept the same as their original paper. We utilized Normal Consistency (NC) and F-Score with thresholds of 0.5% and 1%, denoted as F-0.005 and F-0.01, as the evaluation metrics. We refer the readers to [28] for the detailed definitions. During the evaluation, 500K points were sampled on the deformed and target surfaces. Table I and Fig. 8 show the numerical and visual results, respectively, where both quantitative accuracy and visual quality of deformed shapes with our DDM are much better than baseline methods. While MDA exhibits superior visual results compared to CD and P2F, its numerical results about F-0.005 and F-0.01 in Table I is worse than both. This discrepancy arises from the fact that the deformed surfaces produced by MDA are excessively smoothed, leading to a convergence to local optima, because of its loss function. Although the overall shape is similar, this oversmoothing effect is evident in the error maps illustrated in Fig. 9. Table II presents the running time and GPU memory costs of different distance metrics per iteration. Notably, our DDM incurs a computational cost similar to P2F but higher than CD. This is because both DDM and P2F require identifying the closest points on triangle meshes, a more complex task compared to CD, which only involves finding the nearest points in point clouds.
@@ V-B2 Comparisons
We compared our DDM with EMD [14], CD [15], DCD [17], and ARL [3]. Additionally, we made comparisons with two widely used registration methods: ICP [1] and SICP [48]. We employed two widely used evaluation metrics in the rigid registration task, Rotation Error (RE) and Translation Error (TE), to measure the registration accuracy, respectively defined as
@@
The numerical results are listed in Table III, where the refinement through our DDM further improves the accruacy of MAC significantly, while after the refinement with other distance metric, the registration accuracy even decrease dynamically. Although the RE and TE of SICP are slightly better than ours, its SR is much lower than ours, showing its limitation. The visual results are shown in Fig. 10, and when dealing with the data with noise or outliers, the details of the point clouds become indistinct, our method still register the two point clouds successfully. To evaluate the performance of different methods insightly, we also used the Registration Recall (RR) under different thresholds of RE and TE to measure the registration accuracy. As shown in Fig. 11, the superiority of our DDM is further verified. Additionally, Table IV presents the computational cost per iteration during the optimization process. As shown, EMD incurs the highest computational overhead due to the necessity of constructing a comprehensive bi-directional mapping between two point sets. Similarly, ARL requires significant computational resources, as it involves calculating the intersection of lines with the two point clouds. In contrast, CD, BCD, and our DDM exhibit comparable computational costs. However, DDM achieves superior registration results compared to these methods, demonstrating its effectiveness and efficiency in rigid point cloud registration.
@@ V-C2 Comparisons
We compared our DDM with CD and the P2F distance. For a fair comparison, we kept the number of sampled points in CD and the P2F distance the same as the generated reference points of our DDM. In addition, we also compared with the state-of-the-art optimization-based non-rigid registration method called AMM [4]. We computed RMSE of the estimated and ground truth vertices of the deformed source surfaces for quantitative evaluation. The results in Table V and Fig. 12 demonstrate the significant superiority of our DDM. Additionally, we show the registration errors of all pairs in a sequence in Fig. 13 for a more comprehensive comparison, where it can be seen that the compared methods produce much larger RMSE values for pairs with significant motion, while our DDM always works well. Additionally, Table VI presents the computational time and GPU memory requirements per iteration during the optimization process. Specifically, CD achieves the fastest running speed, as it only involves identifying the nearest points within the point clouds. In contrast, P2F and our proposed DDM require more computational time due to the added complexity of locating the closest point on the triangle mesh, which is inherently more complex than finding the nearest points in a point cloud.
@@ V-D1 Implementation Details
We used the Flyingthings3D dataset [64], where $N_{\rm src}=N_{\rm tgt}=8192$. The hyperparameters involved in our DDM were set as $K=5$ and $M=81920$. In order to adapt the density at each point’s location, the value of $\sigma$ was set as 3 times of the distance to its nearest point. For the optimization-based methods, we optimized the scene flow directly with the Adam optimizer for 500 iterations with a learning rate of 0.01. For unsupervised learning-based methods, we employed two state-of-the-art methods, namely NSFP [12] and SCOOP [13], both of which employ CD as the distance metric. We maintained the training settings identical to those specified in their original papers, with the only modification being the replacement of the alignment criterion with our DDM.
@@ V-D2 Comparisons
Following [12, 13], we employed End Point Error (EPE), Flow Estimation Accuracy (Acc) with thresholds 0.05 and 0.1 (denoted as Acc-0.05 and Acc-0.1), and Outliers as the evaluation metrics. From Table VII and Fig. 14, it can be seen that our DDM drives much more accurate scene flows than EMD, CD, and BCD under the optimization-based framework, and our DDM further boosts the accuracy of SOTA unsupervised learning-based methods to a significant extent, demonstrating its superiority and the importance of the distance metric in 3D data modeling. Table VIII illustrates the computational cost of various distance metrics per iteration during the optimization process. Specifically, EMD demands significantly more computing resources due to the necessity of constructing a bi-directional mapping between two point sets. In contrast, CD, BCD, and our proposed DDM exhibit comparable computational requirements. However, our DDM outperforms the others in terms of flow estimation accuracy, highlighting its superior effectiveness and efficiency.
@@ V-E1 Implementation Details
We used the CAPE dataset [65], which contains numbers of clothed human models and their corresponding SMPL parameters. We selected two sequences from the whole dataset named ‘longlong basketball trial2’ and ‘blazerlong volleyball trial2’. In each sequence, we used the pose of the $i$-th frame as the initial parameter and the points sampled from the ($i$+5)-th frame as the simulated scanned point cloud, where the number of the sampled points was $2\times 10^{4}$. The hyperparameters invovled in our DDM were set as $K=5$, $\beta=1.0$, and $M=6\times 10^{4}$. To bring the experimental setup closer to real-world conditions, where collected point clouds are typically incomplete, we positioned a virtual camera in front of the reference mesh and selectively sampled points exclusively from the visible faces, emulating the data collection process of an RGB-D camera. In such a case, the number of the sampled points was $10^{4}$ and the hyperparameters about our DDM were kept the same. The optimization was conducted through Adam optimizer, running $10^{3}$ iterations with a learning rate of $5\times 10^{-4}$.